Circular No. 5953

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994D IN NGC 4526 R. J. Cumming, W. P. S. Meikle, and T. R. Geballe report: “Our

claimed detection of CO absorption in SN 1994D (cf. IAUC 5951), based on preliminary data reduction by one of us (TRG), is incorrect. More careful reduction (by TRG) of the data demonstrates that the CO absorption is in the underlying galaxy. The same re-reduction reaffirms that the wide absorption feature centered at 1.052 μm is intrinsic to the supernova. In addition, weaker and resolved (but less broad) absorption is apparent at 2.04 μm; this feature is identified as He I 2.058 μm, and it provides strong support for interpretation of the 1.052-μm feature as highly blueshifted He I 1.083-μm absorption.”

Visual magnitude estimates: Mar. 12.37 UT, 12.6 (R. Royer, Lakewood, CA); 13.31, 12.3 (Royer); 14.30, 12.2 (Royer); 15.961, 12.1 (G. M. Hurst, Basingstoke, England); 16.94, 12.2 (G. Comello, Groningen, The Netherlands); 17.399, 11.9 (W. G. Dillon, Missouri City, TX).

GQ MUSCAE M. Hamuy and M. Phillips, Cerro Tololo Interamerican Observatory;

and R. Williams, Space Telescope Science Institute, report: “The optical spectrum of GQ Mus has changed significantly since the last observation in 1992 Jan. Consistent with reports that the x-ray flux has suddenly declined during the past six months (J. Krautter, private communication), the nova appears to be ‘turning off’. The very strong line of [Fe X] 637.5-nm that had been present in the spectrum for the past five years has now disappeared and has been replaced by emission from [Fe VII] 608.7-nm. A spectral scan (range 320–750 nm) obtained with the CTIO 4-m telescope on 1994 Feb. 7 shows that the optical continuum is unchanged, but the emission lines are all substantially weaker and of lower ionization; e.g., [O III] 500.7-nm has once again appeared in the spectrum. This nova should continue to be monitored so that the ‘nova turn-off’ can be documented.”

SUPERNOVA 1982aa IN NGC 6052 The presumed radio supernova in NGC 6052 = Markarian 297 at α =

16^{h}03^{m}00^{s}.99, δ = +20^{o}40′43′′.6 (equinox 1950.0; see Heeschen et al. 1983, Ap.J. 267, L73; IAUC 5276; Yin and Heeschen 1991, Nature 354, 130; Yin 1994, Ap.J. 420, 152) has been given the designation 1982aa.

1994 March 17 Daniel W. E. Green